OpenStudy (cmtboy2):

A recipe calls for 10 cups of oatmeal. Benjamin wants to make fewer servings than the recipe makes. Which expression represents how many cups of oatmeal Benjamin should use? A. 0.8 10 B. 1.8 10 C. 0.2 10 D. 1.2 10

OpenStudy (mathstudent55):

There are numbers and symbols missing in this problem and in the choices.

OpenStudy (cmtboy2):

1/5

OpenStudy (muhammad12):

I cant quite guess what you are saying

OpenStudy (cmtboy2):

1 over 5

OpenStudy (cmtboy2):

like 1/5th

OpenStudy (muhammad12):

I mean about the question you are asking

OpenStudy (mathstudent55):

The recipe calls for 10 cups of oatmeal. He wants to make 1/5 less amount. That means he needs to use 1/5 less oatmeal. I can't understand any of the choices because there are symbols missing.

OpenStudy (cmtboy2):

it's times

OpenStudy (mathstudent55):

Ok. Thanks.

OpenStudy (mathstudent55):

If he is making 1/5 less, that means he is making 1 - 1/5 = 5/5 - 1/5 = 4/5 He is making 4/5 of the recipe. 4/5 as a decimal is 0.8 That means he needs 4/5 or 0.8 of the original amount.

OpenStudy (cmtboy2):

ok

OpenStudy (mathstudent55):

Which choice means 0.8 of the original 10 cups? Remember that in math, "of" usually means multiplication.

OpenStudy (cmtboy2):

ok